When price is equal to 7, a monopolist can sell 30 units of output, and when price is equal to 6, a monopolist can sell 40 units of output.

What is the firm's marginal revenue?
Answer #1

Answer

$3

Total revenue when price is 7 = 7*30=210

Total revenue when price is 6 = 6*40=240

Marginal revenue is the change in revenue due to change in output the formula is:

MR=∆Revenue/∆Output

Or MR = (240-210)/(40-30)

Or MR = $3

(Please consider giving an upvote if you find it useful)